3. (a) If y = a + 13 is a solution of the cubic equation 3
y3= 4y + B = 0 such that 34 = — A, find the
quadratic equation having a 3and 133as
roots.
(b) Without Expanding the determinant 2
1 2 3
4 5 6
7 8 9
find its value.
4. (a) Prove that A u (A n B) = A for any two
subsets A and B of a set U.
(b) Check if the following system of equations
can be solved by cramer's rule. If it can be
solved, then apply the rule for solving the
system ; otherwise solve it by the Gaussian
elimination method.
x+y+z=6
x + 2y + 3z = 10
x -F 2y + 3z = 6.
5. (a) Show that 22n — 3n —1 is divisible by 9 using
the principle of mathematical induction.
(b) Find the greatest value of (3 — x)5(2 + x)6,
for — 2 < x < 3.
1
Expert's answer
2019-11-14T09:54:04-0500

3.(a) Find the quadratic equation having "a=3" and "a=133" as roots.  

Given "a_1=3, a_2=133." Complete the quadratic equation


"C(a-3)(a-133)=0, C\not=0"

"Ca^2 -136Ca+399C=0, C\not=0"


(b)


"\begin{pmatrix}
 1& 2 & 3 \\
 4 & 5 & 6 \\
 7 & 8 & 9
\end{pmatrix}"

Subtract row1 multiplied by 4 from row2 


"\begin{pmatrix}
 1& 2 & 3 \\
 0 & -3 & -6 \\
 7 & 8 & 9
\end{pmatrix}"

Subtract row1 multiplied by 7 from row3 


"\begin{pmatrix}
 1& 2 & 3 \\
 0 & -3 & -6 \\
 0 & -6 & -12
\end{pmatrix}"

Subtract row2 multiplied by 2 from row3


"\begin{pmatrix}
 1& 2 & 3 \\
 0 & -3 & -6 \\
 0 & 0 & 0
\end{pmatrix}"

The determinant


"\begin{vmatrix}
 1& 2 & 3 \\
 4 & 5 & 6 \\
 7 & 8 & 9
\end{vmatrix}=\begin{vmatrix}
 1& 2 & 3 \\
 0 & -3 & -6 \\
 0 & 0 & 0
\end{vmatrix}=0"

4. (a)


"A\cap B=\{{x:x\in A\ and\ x\in B}\}"

"A\cup(A\cap B)=\{{x:x\in A\ or ( \ x\in A\ and\ x\in B)}\}"

Hence


"A\cup(A\cap B)=\{{x:x\in A}\}=A"



(b)


"x+y+z=6""x+2y+3z=10""x-2y+3z=6"

"\varDelta=\begin{vmatrix}
 1& 1 & 1 \\
 1 & 2& 3 \\
 1 & -2& 3
\end{vmatrix}=""=1\begin{vmatrix}
 2 & 3 \\
 -2 & 3
\end{vmatrix}-1\begin{vmatrix}
 1 & 3 \\
 1 & 3
\end{vmatrix}+1\begin{vmatrix}
 1 & 2 \\
 1 & -2
\end{vmatrix}=""=(2(3)-3(-2))-(1(3)-3(1))+(1(-2)-2(1))=""=12-0-4=8\not=0"

"\varDelta_x=\begin{vmatrix}
 6 & 1 & 1 \\
 10 & 2& 3 \\
 6 & -2 & 3
\end{vmatrix}=""=6\begin{vmatrix}
 2 & 3 \\
 -2 & 3
\end{vmatrix}-1\begin{vmatrix}
 10 & 3 \\
 6 & 3
\end{vmatrix}+1\begin{vmatrix}
 10 & 2 \\
 6 & -2
\end{vmatrix}=""=6(2(3)-3(-2))-(10(3)-3(6))+(10(-2)-2(6))=""=72-12-32=28"

"\varDelta_y=\begin{vmatrix}
 1& 6 & 1 \\
 1 & 10 & 3 \\
 1 & 6 & 3
\end{vmatrix}=""=1\begin{vmatrix}
 10 & 3 \\
 6 & 3
\end{vmatrix}-6\begin{vmatrix}
 1 & 3 \\
 3 & 3
\end{vmatrix}+1\begin{vmatrix}
 1 & 10 \\
 1 & 6
\end{vmatrix}=""=(10(3)-3(6))-6(1(3)-3(1))+(1(6)-10(1))=""=12-0-4=8"

"\varDelta_z=\begin{vmatrix}
 1& 1 & 6 \\
 1 & 2& 10 \\
 1 & -2 & 6
\end{vmatrix}=""=1\begin{vmatrix}
 2 & 10 \\
 -2 & 6
\end{vmatrix}-1\begin{vmatrix}
 1 & 10 \\
 1 & 6
\end{vmatrix}+6\begin{vmatrix}
 1 & 2 \\
 1 & -2
\end{vmatrix}=""=(2(6)-10(-2))-(1(6)-10(1))+6(1(-2)-2(1))=""=32+4-24=12"

"x={\varDelta_x \over \varDelta}={28 \over 8}={7 \over 2}"

"y={\varDelta_y \over \varDelta}={8 \over 8}=1"

"z={\varDelta_z \over \varDelta}={12 \over 8}={3 \over 2}"

"({7 \over 2},1, {3 \over 2})"

5. (a)   

 Let the statement "P(n)" given as


"P(n):2^{2n}-3n-1\ \text{is divisible by 9,}n=1, 2, ...."

Base case:

"n=1: P(1)=2^{2(1)}-3(1)-1=0," is divisible by "9."

Assume that "P(n)" is true for some natural number "k," i.e.,


"P(k):2^{2k}-3k-1\ \text{is divisible by 9,}\text{i.e.}""2^{2k}-3k-1=9q, q\in \N"

Now, to prove that "P(k+1)" is true, we have


"P(k+1)=2^{2(k+1)}-3(k+1)-1=""=4\cdot2^{2k}-3k-3-1=""=4\cdot2^{2k}-12k-4+12k+4-3k-3-1=""=4\cdot(2^{2k}-3k-1)+9k=4(9q)+9=""=9(4q+1)=9m, m\in \N"

Thus "P(k+1)" is true, whenever "P(k)" is true. Hence, by the Principle of Mathematical Induction "P(n)" is true for all natural numbers "n."


(b)


"f(x)=(3-x)^5(2+x)^6,\ -2<x<3"

"f'(x)=-5(3-x)^4(2+x)^6+6(3-x)^5(2+x)^5="

"=(3-x)^4(2+x)^5(-10-5x+18-6x)="

"=(3-x)^4(2+x)^5(8-11x)""f'(x)=0=>(3-x)^4(2+x)^5(8-11x)=0"

Find critical number(s)


"-2, {8 \over 11} , 3"

First Derivative Test


"\text{If }-2<x<{8 \over 11}, f'(x)>0, f\text{ increases}""\text{If }{8 \over 11}<x<3, f'(x)<0, f\text{decreases}"

The greatest value of "(3-x)^5(2+x)^6" for "-2<x<3"

"f({8 \over 11})=(3-{8 \over 11})^5(2+{8 \over 11})^6={25^5\cdot30^6 \over 11^{11}}\approx"

"\approx24952"